OpenStudy (anonymous):

Let f(x) = e^x - e^(3 x). Find all extreme values (if any) of f on the interval 0 <= x <= 1. Determine at which numbers in the interval these values occur. Remember to check for endpoint extrema

OpenStudy (anonymous):

how do i even start?

OpenStudy (anonymous):

after finding the deriv of course

OpenStudy (anonymous):

f'(x)= e^x-3e^3x

sam (.sam.):

I don't think there's any

sam (.sam.):

In interval 0 <= x <= 1

OpenStudy (anonymous):

no max values but there are min values

sam (.sam.):

You sure there are min values?

sam (.sam.):

If you're talking from -1<=x<=0 Then we'll have maximum

OpenStudy (anonymous):

im this online math study site and the problem has answers and it has only a min value of 1

OpenStudy (anonymous):

max/min values occur at the endpoints on the interval [0, 1]

sam (.sam.):

That's weird @satellite73 @Hero @TuringTest @bahrom7893

sam (.sam.):

I have to go

OpenStudy (bahrom7893):

f'(x)=e^x-3e^(3x)

OpenStudy (bahrom7893):

e^x(1-3e^(2x))=0 1-3e^(2x)=0 e^(2x)=1/3 2x=-ln3 x=-(ln3)/2

OpenStudy (anonymous):

your problem asks for extreme values on [0, 1]... according to the graph it is only decreasing on this interval. so the max must occur at x=0 and min at x=1.

OpenStudy (bahrom7893):

ohh correct dpalnc, i didn't even look at the graph nor read the question properly.

OpenStudy (anonymous):

youre right dpalnc

OpenStudy (anonymous):

i got that already but what is the crit value for the min ?

OpenStudy (anonymous):

if you want the value of the min in [0, 1], it's f(1)...

OpenStudy (anonymous):

ok thanks